The inductance of a solenoid with 420 turns and a length of 21cm is 7.6mH .

What is the cross-sectional area of the solenoid?

What is the induced emf in the solenoid if its current drops from 3.0A to 0 in 60ms ?
Answer #1

(a)
L = uoN2A / l
7.6 x 10-3 = 4*pi*10-7 x 4202 * A / 0.21
A = 7.2 x 10-3 m2

(b)
EMF, E = -Nd(flux)/dt
but L = Nd(flux) / dI
Nd(flux) = L*dI
So,
E = -L dI/dt
E = -7.6 x 10-3 * (0 - 3) / 60 x 10-3
E = 0.38 V